Find the fifth, ninth and hundredth term of the sequence. Show your work for full credit!
4, 14, 24, 34, ...
A(n) = 4 + (n - 1)(10)

Answer by ewatrrr(24785)

4, 14, 24, 34, ...
A(n) = 4 + (n - 1)(10) Good work, d = 10
.......
A(5) = 4 + (n - 1)(10) = 4 + 4(10) = 44
A(9) = 4 + (n - 1)(10) = 4 + 8(10) = 84
A(100) = 4 + (n - 1)(10) = 4 + 99(10) = 994